OpenStudy (anonymous):

If y = 3^x(lnx^2) , find dy/dx.

OpenStudy (anonymous):

Did you just need help calculating it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Is the equation? \[3^{x}*\ln(x^2)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We would need to apply both the prodcut rule and chain rule

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

We would need to start with the product rule so we take the derivative of the (first term times the second term without doing the derivative of it) plus (the first term times the derivative of the second term)

OpenStudy (anonymous):

there is a quick little formula for finding the derivative of a constant to the x power

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

If we have a to the x (where a is a constant)....if we want to find the derivative of it we would get (a^x)lna

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so go ahead and trying doing it

OpenStudy (anonymous):

okay i'll try thanks

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i got dy/dx = 2(3^x)/x - 3^x(ln3)(lnx^2)

OpenStudy (anonymous):

let me quickly calculate it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[3^x*3\ln(x^2)+(2x/x^2)*3^x\]

OpenStudy (anonymous):

Thats what I got and I'm pretty sure it is right

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Sorry I need to go but good luck